OpenStudy (anonymous):

Part 1. Identify p, q such that x+5x+6=(x+p)(x+q) Part 2. Find the roots of f(x)=x+5x+6

OpenStudy (anonymous):

p=−x2+xq−6x−6 over q+x and q=−x2+xp−6x−6 over p+x this is part 1

OpenStudy (anonymous):

x=-1 is the root for part 2